# General task optimal planning for heterogeneous teams with precedence and compatibility constraints and its application on power grid inspection using unmanned aerial vehicles

### v1

@@ Abstract
In this paper we address the optimal planification of general purpose tasks that includes a wide spectrum of situations: from project management of human teams to the coordination of an automated assembly line or the automated inspection of power grids.
There exists many methods for planification. However, the vast majority of such methods are conceived for very specific problems or situations. The main consequences of this is that no general planification method exists and the rigidity that prevents the extension of current methods to new cases and applications.
To address this, we propose a new truly general method ultimately based on the generalization of the Travelling Salesman Problem (TSP) that we call the Heterogeneous Multiworker Task Planification Problem (HMWTPP). The HMWTPP is then used to model and solve several classical problems included in the TSPLIB [1] library for validation. We then solve an example of an assembly line to show the capabilities and flexibility of the HMWTPP.
To conclude, we adapt the HMWTPP to the planification of unmanned aerial vehicles (UAVs), specifically to the automated inspection of power grids. This adaptation was validated by solving real-life cases for power grids in ATLAS Flight Test Center at Villacarrillo, Spain.
@@ 1 Introduction
In modern times, due to the exponential increase in the number of workers involved, coordination between members of a group or team, human or robot, is key. This is even more important as entire industries are moving towards automation due to the advancements in robotics and artificial intelligence.
@@
As of today, task planification is either made manually or using mathematical models based on the TSP problem. The issue with these methods is that they are either suboptimal or lack flexibility, making them hard to use in problems for which they were not designed. Although many of them are used [2, 3, 4, 5, 6, 7, 8] and cover very diverse applications (vehicle routing, transportation and delivery, network connectivity, search and rescue and agriculture to name a few). However, no mathematical formulation brings them all under the same umbrella.
@@
In this paper, we present a general task optimal planification model based on a generalization of the Traveling Salesman Problem (TSP) called the Multiworker Task Planification Problem (HMWTPP) [^1] Which is closely related to the Multi-Robot Task Allocation Problem (MRTAP). It uses the abstract concept of task and worker to model generic planification problems that are not only useful for robotics but in many generic situations where a team is required to complete a set of tasks. This model is presented in section 2, where all the mathematical foundation is established. We show that an instance of the model can be represented by a directional multigraph in conjunction with a Mixed Integer Linear Problem (MILP) solvable by any off-the-shelf optimization suites. We then use it to solve several classical problems from the TSPLIB [1] in subsection 2.3 and the planification of a guitar assembly line in subsection 2.4. This last example includes many of the unique quirks of our model and shows its versatility.
@@
To conclude, we present a more complex problem: the inspection of power grids using UAVs in section 3. We describe the process of the implementation of UAVs in such scenario and select a specific emplacement, the ATLAS Flight Test Center for real-world planification to further test and validate the HMWTPP model. The solution for several instances of problem for the ATLAS power grid are discussed.
2 The HMWTPP formulation for general task optimal planification problem
@@
As we will see, the HMWTPP can be described using a weighed directed multigraph coupled with a mixed-integer linear problem (MILP). The corresponding MILP is an optimization problem that encodes all the information needed and which solution exactly solves the planification problem. To find such solution, either any off-the-shelves optimization suites or custom branch-and-cut algorithms can be used without much modification to the code or the solver.
@@ 2.1 Weighed directed multigraph representation
Consider a weighted directed multigraph $G=(V,E,W)$ where $V$ is the set of vertices, $E$ is the set of edges and $W$ the set of weights, and a team of heterogeneous workers $\mathcal{W}$.
The set of workers $\mathcal{W}$ contains all the information needed to characterize and mathematically model each of the workers $w\in\mathcal{W}$. This team is required to complete a list of tasks $\mathcal{T}$ in a cooperative and optimal manner. Each of the tasks $\tau\in\mathcal{T}$ can be performed in $n_{\tau}$ different approaches ${\mathcal{T}_{\tau}=\{\tau^{(n)}\}_{n=1}^{n_{\tau}}}$, i.e., the different ways a task can be carried out. Each of the elements in the set of all possible task approaches $\hat{\mathcal{T}}=\cup_{\tau\in\mathcal{T}}\mathcal{T}_{\tau}$ is associated with a vertex $v_{\tau}^{(n)}\in V$. The set of all the vertices associated with an specific task is $T_{\tau}\subset V$ and the set containing all of such vertices is $T=\cup_{\tau\in\mathcal{T}}T_{\tau}\subset V$, called the set of tasked vertices.
To allow workers to start and finish their route we define the bases set $B\subseteq V$ (${B\cap T=\emptyset}$). Bases can either represent physical points or the abstract command to start and finish the route. The number of bases $b\in B$ must be at least one and has no upper bound. All workers $w$ need to have exactly one base $b_{w}$ assigned to them although one can be shared by several workers. ($b_{w_{1}}=b_{w_{2}}$ for some $w_{1},w_{2}\in\mathcal{W}$).
Due to the possible heterogeneity of the workers, some of the task approaches might be only compatible with a subset of the workers. To easily express this mathematically in the model we define, for any subset $S$ of $\hat{\mathcal{T}}$ (and $T$ by extension), the $w$-restriction operation $\cdot\,|_{w}$ that selects the subset of $S$ compatible with the worker $w$.
$S|_{w}=\left\{s\in S\,|\,\Gamma_{w}(s)=1\right\}$ (1)
@@
This allow us to decompose the set of tasks (and tasked vertices) into compatibility subsets:
$\hat{\mathcal{T}}=\bigcup_{\forall w\in\mathcal{W}}\hat{\mathcal{T}}\lvert_{w};\quad T=\bigcup_{\forall w\in\mathcal{W}}T\lvert_{w}=\bigcup_{\forall w\in\mathcal{W}}\bigcup_{\forall\tau\in\hat{\mathcal{T}}\lvert_{w}}v_{\tau}$ (3)
Notice that for any pair of workers $w_{1},w_{2}\in\mathcal{W}$, $\hat{\mathcal{T}}\lvert_{w_{1}}\cap\,\hat{\mathcal{T}}\lvert_{w_{2}}\neq\emptyset$ ($T\lvert_{w_{1}}\cap\,T\lvert_{w_{2}}\neq\emptyset$) in general as a task can be compatible with more than one worker.
To represent the transition of the worker $w$, from the task approach $\tau_{1}^{(a)}$ to the task approach $\tau_{2}^{(b)}$, we use the edges $E$ of $G$. One edge $e\equiv[v_{\tau_{1}}^{(a)},v_{\tau_{2}}^{(b)}]|w$ is labeled by an ordered pair of vertices $[v_{\tau_{1}}^{(a)},v_{\tau_{2}}^{(b)}]$ and a worker $w$ (which makes up for the multi in the name multigraph). In a sense, this multigraph can be seen as a layered set of regular graphs, one for each worker, coupled by the common vertices.
The set $E$ must encode all the information about task compatibility and possible task transitions. For this reason not all possible connections between vertices are allowed. Iterating over each of the workers $w$, the assigned base $b_{w}$ (remember that some workers may share the same) is connected to each of the vertices of $T|_{w}$ in both directions. Then, all the vertices from $T|_{w}$ are connected with each other except for those that are related to the same task, i.e., the elements of any of the sets $T_{\tau}$. This prevents the completion of the same task twice and avoids loops. Additionally, some tasks may require extra rules for the its connections. Consider for example the case where a worker builds a house roof before its foundations. An advantage of imposing the compatibility directly into the connections of the graphs versus doing so as external constraints is that the number of edges decreases and therefore the complexity of $G$. As we will see, this directly translate into a simpler problem to solve. Considering the case where all tasks can be done in $n_{a}$ different approaches by all workers, the number of edges in the worst case scenario is $\propto n_{w}(n_{a}n)^{2}$ (where $n$ is the number of tasks and $n_{w}$ is the number of workers) which rapidly scales into a numerical nightmare if the problem is too complex and is not solved carefully.
To quantify the costs of completing a task or transitioning between them, each of the edges $e\equiv[u,v]|w\in E$ is provided with a set of weights $\{\Omega^{(\mu)}_{e}\}_{\mu,e\in E}$. They include the cost associated with the transition from $u$ to $v$ and the execution of the task at $v$ for the worker $w$.
$\Omega^{(\mu)}_{[u,v]|w}=\Delta\omega^{(\mu)}_{[u,v]|w}+\omega^{(\mu)}_{v|w}$ (4)
@@
The weights $W$ needs to be computed beforehand so their value does not explicitly depends on the worker’s route. One way to avoid this limitation is to check whether a certain condition or events happens mid-route and re-plan the routes once this happens.
@@ 2.2 Mixed-integer programming problem (MILP) representation
Once the weighted directed multigraph $G$ is fully constructed and all the external information of the problem is encoded in it, we would still lack a description of the actual mathematical problem and its solution. Consider one of the workers, How is its route, that is the order at which it performs the assigned tasks, described in the context of $G$? It is represented as a path $p$ in $G$, an unordered list of adjacent edges (which considering the visited vertices, is also a subgraph of $G$). Not at all paths are valid solutions however. Some extra criteria and constraints are still needed [^2] Different starting and ending points can be easily defined if one ignores the last edge of an already planed route and manually adds the transition to the end point. This way, internally the route still closes back to the base which simplifies some aspects of the model.. For instance, all valid routes must start and finish at the corresponding base, so all routes that do not fulfil this are automatically discarded.
To numerically represent a path $p_{w}$, we define a set of binary parameters $\mathbf{Z}=\left\{z_{e}\right\}_{e\in E}$ that activates certain edges of $G$ in a route:
@@
To ensure, however, that a given choice of $\mathbf{Z}$ represents a set of valid routes, it needs to fulfil some additional conditions. Let’s first introduce some useful notation in a similar fashion to that of [5].
@@
(notice that the double sum might yield non-existent connections, however, those terms add no contribution to the sum). These operators measure how many edges are active and exiting ($+$) or entering (--) $S$ for the worker $w$, i.e, the number of times the worker enters and exits the subset. We also define the absolute divergence operator $\nabla^{(w)}=\nabla^{(w)}_{+}+\nabla^{(w)}_{-}$. By abuse of notation, we can use any of these operators over a vertex $v\in V$ using the same notation used for sets.
@@
is also needed. It represents the total number of connections within a subset of $V$ for a worker.
For a given $u\in V$ and a worker $w$ we define the following sets:
@@
With this notation, we can easily formulate the conditions or constraints needed to define valid paths.
To ensure that all workers $w\in\mathcal{W}$ start and finish their route at their assigned base $b_{w}\in B\subset V$ we use the bases constraints ($\mathcal{C}_{\text{B}}$) Equation 11. To allow some of the worker to remain inactive, we introduce some new binary free parameters $\mathbf{Y}_{B}=\left\{y_{w}\in\{0,1\}\right\}_{w\in\mathcal{W}}$
@@
Usually, it is simpler to set $y_{w}=1$ for $\forall w\in\mathcal{W}$ and simplify all the rest of the problem by eliminating any of the references to $\mathbf{Y}_{B}$. This however might yield sub-optimal solutions or even create contradiction if there are less tasks than workers. As previously said, if we wanted to make a worker start and finish at two different locations, we could just manually delete the last edge on its route after planification and consider the last transition as the one to the new end point. The planification process remains the exact same and internally the route still closes back to itself on the base.
As all task must be completed by a compatible (remember that compatibility is encoded in the edges of the graph) and active worker, exactly once in any of the available approaches, we impose the task completion constraints ($\mathcal{C}_{\text{T}}$) Equation 12 and Equation 13. We also introduce the parameters ${\mathbf{Y}_{C}=\{y_{v|w}\in\{0,1\}\}_{w\in\mathcal{W};\,v\in T}}$ to select which worker is to complete each of the tasks and define ${\mathbf{Y}=\mathbf{Y}_{C}\cup\mathbf{Y}_{B}}$. In a similar fashion to that of $\mathbf{Z}$, they are separated into fixed and free parameters. If a task $\tau$ is incompatible with $w$ ($\Gamma_{w}(\tau)=0$), then, $y_{v_{\tau}|w}$ is fixed at 00. All the parameters that correspond to compatible associations, are left free. We then need to add the following constraints:
@@
They also enforce continuity between transitions.
Still, it is possible that the routes yielded by a specific selection of $\mathbf{Z}$ and ${\mathbf{Y}=\mathbf{Y}_{B}\cup\mathbf{Y}_{C}}$ contains several closed loops that do not share any vertex called subtours. This an unwanted side effect that we need to remove from the model. To do so, we need to add some kind subtour elimination constraints ($\mathcal{C}_{\text{S}}$). In literature, two types are mainly used: the Dantzig–Fulkerson–Johnson [9] (DFJ SECs) or the Miller-Tucker-Zemlim [10] subtour elimination constraints (MTZ SECs). Each of them have their own unique pros and cons.
@@
$\left\{\Sigma^{(w)}S\leq|S|-1\right\}_{\begin{subarray}{c}W\in\mathcal{W}\\ S\in\{Q\subset V\,|\,b_{u}\notin Q,\,|Q|\geq 2\}\end{subarray}},$ (14)
and each of them enforces that for $Q\subset V$, the number of moves done by worker within the subset is exactly the minimum number of moves needed to transverse all of its vertices. This way, no valid route can loop to itself within $Q$. As we know where each of the worker starts and finishes, we do not need to add the condition for all the possible subsets of $V$, just for those that do not contain the base.
The main advantage of these constraints is their simplicity and generality that allow for a more dynamic solving process without any modification to their original constraint’s formulation. The number of subtour constraints scale exponentially as so does the number of possible subsets $Q$. For instances where the number of tasked vertices is significant, the solving process with all the necessary constraints may take an absurd amount of computation effort so a different approach is needed. Instead of solving the complete problem, we relax it by removing the DFJ constraints. We then solve this relaxed instance and check for subtours. If subtours are found, we can eliminate them by adding the corresponding DFJ from Equation 14 for all $Q$ that are a subset of the set of vertices within each of the invalid subtours and then iterate again in a branch-and-cut fashion. In many cases, this approach speeds thing up by orders of magnitude depending on the specific problem and alleviate some of the computing requirements. However, this heavily depends on the number of iterations one needs to make until the true optimal solution is found. It is also very limited in terms of extensions to the formulation.
On the other hand, although the process is not straightforward the MTZ constraints can be easily adapted to this formulation by modifying the originals to:
$\left\{p_{u|w}-p_{v|w}+1\leq m_{w}(1-z_{[u,v]|w})\right\}_{\begin{subarray}{c}w\in\mathcal{W}\\ u,v\in T\lvert_{w}\end{subarray}},$ (15)
and introducing two auxiliary constraints:
@@
$\left\{z_{[b_{w},u]|w}\leq p_{u|w}\right\}_{\begin{subarray}{c}w\in\mathcal{W}\\ u\in\delta^{(w)}_{+}(b_{w})\end{subarray}}$ (17)
where $m_{w}=|T\lvert_{w}|$ and $\mathbf{P}=\left\{p_{u|w}\in\mathbb{N}\right\}_{w\in\mathcal{W};\,u\in T\lvert_{w}}$ are new free parameters that encode the order at which a vertex $u\in T\lvert_{w}$ is visited within the route of $w\in\mathcal{W}$. The constraints Equation 15 enforces that if $u$ is visited before $v$ by $w$, then $p_{u|w}\geq p_{v|w}$. Notice however that if $v$ is visited immediately after $v$, $p_{v|w}$ might not be equal to $p_{u|w}+1$ as the number of vertices visited within a route might be less than $m_{w}$ so Equation 16 is trivially fulfilled. At the same time, Equation 16 and Equation 17 assigns $p=0$ for non-visited vertices and at least $p=1$ to the first one in each of the routes respectively. These constraints can be easily tightened so the values of $\mathbf{P}$ directly match with the visit order by defining $m_{w}=\sum_{e\in E\lvert_{w}}z_{e}$ and linearizing the products of the type $z_{e}m_{w}$. This however is not very useful and much of the actual useful characteristic of the MTZ SECs are also available in this formulation.
As the original MTZ SECs, they provides a natural framework for visit order constraints, however, they can not be used for an relax-solve-iterate solving approach as they all must be included simultaneously to function properly.
Depending on the problem and if the MTZ SECs are used, it might be useful to add constraints into the order of visit of each vertex in a route. For example, consider that the worker $w$ must visit $u$ before $v$ (denoted as $u<_{w}v$) for some technical reason. To implement this behaviour, we would need to impose a new type of constraint in the model, the order of visit constraints $\mathcal{C}_{\text{O}}$. For each of the workers $w$ and pair of order-constrained vertices $[u,v]$, the following constraint must be added:

### v2

# Optimal planning for heterogeneous autonomous teams with precedence and compatibility constraints and its application on power grid inspection with Unmanned Aerial Vehicles

@@ Abstract
In this paper we address the optimal planning of autonomous teams for general purpose tasks including a wide spectrum of situations: from project management of human teams to the coordination of an automated assembly lines, focusing in the automated inspection of power grids.
There exist many methods for task planning. However, the vast majority of such methods are conceived for very specific problems or situations and are often based in certain assumptions and simplifications. Consider for example all the different algorithms developed to solve the Vehicle Routing Problem (VRP) for all the different vehicles and environment characteristics. This means that no robust general planning method exists and that a possible extension of any of them to a more general situation is often not a trivial task.
To address this, we propose a new truly general method ultimately based on a generalization of the Traveling Salesman Problem (TSP). We call this new model the Heterogeneous Multi-worker Task Planning Problem (HMWTPP). It provides a natural framework to model many situations typical in task planning of all kinds. Task-Worker compatibility, precedence/order and time-windows constraints are already encoded into the HMWTPP while it can be easily extended to include weight capacity or battery per node constraints in an intuitive manner.
Several classical TSP problems included in the TSPLIB library are solved for validation and performance analysis of HMWTPP showing a comparable numerical performance to that of existing models. In addition, a synthetic example modeling an automated assembly line is analyzed to prove the potential capabilities of the HMWTPP in real-life scenarios.
Ultimately, we focus in the computation of the optimal plan of Unmanned Aerial Vehicles (UAVs) specifically in the context of automated inspection of electrical power grids. This adaptation is validated by solving real-life missions for a sample power grid segment in the ATLAS Flight Test Center at Villacarrillo, Spain.
@@ 1 Introduction
In modern times, due to the exponential increase in the number of workers involved, coordination between members of a group or team, human or automata, is key. This is even more important as entire industries are moving towards complete automation due to the advancements in robotics and artificial intelligence. As of today, task planning, at least in the context of Unmanned Aerial Vehicles (UAVs), is either made manually or computed numerically using mathematical models often based on the Traveling Salesman Problem [1]. The issue with many of these models is that they are either too specific or lack flexibility, making them a not so good choice for problems for which they were not originally designed. Although many of them are have been proved to work efficiently [2, 3, 4, 5, 6, 7, 8] and cover very diverse applications (vehicle routing, transportation and delivery, network connectivity, search and rescue and agriculture to name a few), they are disseminated and no unique model brings all the different aspects of these models into the same umbrella.
In general, many task planning problems can be described using concepts common to all of them. A series of tasks, which can be either physical or abstract, needs to be performed by a single human or automata or team of them in an optimal manner based on a certain metric. In addition, additional constraints can be imposed. Although their classification is difficult, two main categories arise:
  • Per worker constraints imposed only to a specific worker. They include weight capacity constraints in the context of package delivery, battery level in the context of UAVs planning or task compatibility constraints. As an example, the same mathematical expressions for this type of constraint can also be used to model worker fatigue levels in a project management context.
  • Global constraints imposed over the entire team. They include time windows constraints, i.e., constraints over when a task needs to be performed. They also include constraints that forbid specific feasible solutions.
As examples, in [9] although weight capacity, energy and time windows constraints are fully modeled, only a single worker is considered. In [10], although an entire team is considered, it is not energy-aware and the characteristics of the already chosen team members are directly encoded into the model making it impossible to add compatibility constraints. In [11], authors propose a TSP-based model for UAV path planning that it is not only energy-aware but also compatible with autonomous recharging. It is however limited to a single UAV and does not contain precedence/order nor task compatibility constraints.
@@
In this paper, we present a general task optimal planning model based on a generalization of the Traveling Salesman Problem (TSP) called the Multiworker Task Planning Problem (HMWTPP) [^2] Which is closely related to the Multi-Robot Task Allocation Problem (MRTAP) that brings many of the aforementioned characteristics into a single mathematical framework. It uses the abstract concept of task and worker to model generic planning problems that are not only useful for robotics but in many generic situations where an autonomous team is required to complete a set of tasks. This model is presented in section 2, where all the mathematical foundation is established. We show that an instance of the model can be represented by a directional multigraph in conjunction with a Mixed Integer Linear Problem (MILP) solvable by any off-the-shelf optimization suites. We then use it to solve several classical problems from the TSPLIB [12] in subsection 2.3 and the planning of a guitar assembly line in subsection 2.4. This last example includes many of the unique quirks of our model and shows its versatility.
@@
To conclude, we present a more complex problem: the inspection of power grids using UAVs in section 3. We describe the process of the implementation of UAVs in such scenario and select a specific emplacement, the ATLAS Flight Test Center for real-world planning to further test and validate the HMWTPP model. The solution for several instances of problem for the ATLAS power grid are discussed.
2 The HMWTPP formulation for optimal task planning problems
@@
As we will see, the HMWTPP can be described using a weighed directed multigraph (WDM Graph) coupled with a mixed-integer linear problem (MILP). The corresponding MILP is an optimization problem that encodes all the information needed to completely describe the HMWTPP numerically and to compute its solution. Any off-the-shelf optimization solver can be used to find it.
@@ 2.1 Weighed directed multigraph representation
Consider a weighted directed multigraph $G=(V,E,W)$ where $V$ are the vertices, $E$ are the edges and $W$ are the weights of $G$, and a team of heterogeneous workers $\mathcal{W}$.
The set $\mathcal{W}$ contains all the information needed to characterize and mathematically model each of the team workers $w\in\mathcal{W}$. This team is required to complete a list of tasks $\mathcal{T}$ in a cooperative and optimal manner. Each of the tasks $\tau\in\mathcal{T}$ can be performed in $n_{\tau}$ different approaches ${\mathcal{T}_{\tau}=\{\tau^{(n)}\}_{n=1}^{n_{\tau}}}$, i.e., the different ways a task can be carried out. Each task approach in the set $\hat{\mathcal{T}}=\cup_{\tau\in\mathcal{T}}\mathcal{T}_{\tau}$ is associated with a vertex $v_{\tau}^{(n)}\in V$. The set of all the vertices associated with an specific task is $T_{\tau}\subset V$ and the set containing all of such vertices is $T=\cup_{\tau\in\mathcal{T}}T_{\tau}\subset V$, called the set of tasked vertices.
To allow the workers to start and finish their route we define the set of vertices $B\subseteq V$ (${B\cap T=\emptyset}$) or the bases. Bases can either represent a physical deployment point or the abstract command to start and finish the route. The number of bases $b\in B$ must be at least one and all workers $w\in\mathcal{W}$ need to have exactly one base $b_{w}$ assigned to them although one base can be shared by several ($b_{w_{1}}=b_{w_{2}}$ for some $w_{1},w_{2}\in\mathcal{W}$).
Due to the possible heterogeneity of the team, some of the task approaches might be only compatible with a subset of the workers. To easily express this mathematically, we define, for any subset $S$ of $\hat{\mathcal{T}}$ (and $T$ by extension), the $w$-compatible restriction operation $\cdot\,|_{w}$ that selects the subset of $S$ that is compatible with the worker $w$:
$S|_{w}=\left\{s\in S\,|\,\Gamma_{w}(s)=1\right\},$ (1)
@@
This allows us to decompose the set of tasks (and tasked vertices) into compatibility subsets:
$\hat{\mathcal{T}}=\bigcup_{\forall w\in\mathcal{W}}\hat{\mathcal{T}}\lvert_{w};\quad T=\bigcup_{\forall w\in\mathcal{W}}T\lvert_{w}$ (3)
Notice that in general for any pair of workers $w_{1},w_{2}\in\mathcal{W}$, ${\hat{\mathcal{T}}\lvert_{w_{1}}\cap\,\hat{\mathcal{T}}\lvert_{w_{2}}\neq\emptyset}$ ($T\lvert_{w_{1}}\cap\,T\lvert_{w_{2}}\neq\emptyset$) as a task might be compatible with more than one worker.
To represent the transition of the worker $w$ from the task approach $\tau_{1}^{(a)}$ to the task approach $\tau_{2}^{(b)}$, we use the edges $E$ of $G$. One edge $e\equiv[v_{\tau_{1}}^{(a)},v_{\tau_{2}}^{(b)}]|w$ is labeled by an ordered pair of vertices $[v_{\tau_{1}}^{(a)},v_{\tau_{2}}^{(b)}]$ and a worker $w$ (which makes up for the multi in multigraph). In a sense, this multigraph can be seen as a layered set of regular graphs, one for each worker, coupled by common vertices.
The set $E$ must encode all the information about task compatibility and possible task transitions. For this reason, not all possible connections between vertices are allowed. Iterating over each of the workers $w$, its base $b_{w}$ (remember that some workers may share the same) is connected to each of the vertices of $T|_{w}$ in both directions. Then, all the vertices from $T|_{w}$ are connected with each other except for those that are related to the same task, i.e., vertices from each of sets $T_{\tau}$ cannot be connected with vertices from the same set. This prevents the execution of the same task twice and avoids trivial loops in the graph. Additionally, some tasks may require extra rules for its connections.
An advantage of imposing the compatibility directly into the connections of the graphs versus doing so using external constraints is that the number of edges decreases and therefore the complexity of $G$. As we will see, this directly translates into a simpler problem to solve and a reduced computational time. Considering the case where all tasks can be done in $n_{a}$ different approaches by all workers, the number of edges in the worst-case scenario is $\propto n_{w}(n_{a}n)^{2}$ (where $n$ is the number of tasks and $n_{w}$ is the number of workers) which rapidly scales into a numerical nightmare if the problem is too complex and is not solved carefully. Although no proof is given, it is easy to see that this problem is at least NP-Hard as it is a more complex version of the TSP [13].
To quantify the costs of completing a task and the transitions between them, each of the edges $e\equiv[u,v]|w\in E$ is provided with a set of weights $\{\Omega^{(\mu)}_{e}\}_{\mu}\in W$. They include the cost associated with the transition from $u$ to $v$ and the execution of the task at $v$ for the worker $w$.
$\Omega^{(\mu)}_{[u,v]|w}=\Delta\omega^{(\mu)}_{[u,v]|w}+\omega^{(\mu)}_{v|w},$ (4)
@@
The weights $W$ needs to be computed beforehand so their value does not explicitly depend on the worker’s path. One way to avoid this limitation is to check whether a certain condition or events happens mid-route and re-plan the routes with such route forbidden as a solution. This is shown in subsection 2.2.
@@ 2.2 Mixed-integer programming problem (MILP) representation
Although the WDM graph $G$ is fully constructed, we still lack a description of the actual mathematical problem and its solution. Consider one of the workers, how is its route, i.e., the order at which it performs the assigned tasks, described in the context of $G$? It is represented as a cycle $c$ in $G$, an unordered list of unrepeated adjacent edges (which considering the visited vertices, is also a subgraph of $G$) that loops back to its start vertex. Not at all cycles are valid solutions however. Some extra criteria and constraints are still needed. For instance, all valid routes must start and finish at the corresponding base [^3] Different starting and ending points can be easily defined if one ignores the last edge of an already planed route and manually adds the transition to the end point. This way, internally the route still closes back to the base which simplifies some aspects of the model., so all routes that do not fulfill this are automatically discarded.
To numerically represent a cycle $c_{w}$, we define a set of binary parameters $\mathbf{Z}=\left\{z_{e}\right\}_{e\in E}$ that activates certain edges of $G$ in a route:
@@
As already mentioned, to ensure that a given choice of $\mathbf{Z}$ represents a set of valid routes, it needs to fulfill some additional conditions. Let’s first introduce some useful notation in a similar fashion to that of [5].
@@
(notice that the double sum might include non-existent edges, however, those terms yield no contribution). These operators measure how many active edges related to the worker $w$ are exiting ($+$) or entering (--), i.e, the number of times the worker enters and exits the subset. We also define the absolute divergence operator $\nabla^{(w)}=\nabla^{(w)}_{+}+\nabla^{(w)}_{-}$. By abuse of notation, we can use any of these operators over a single vertex $v\in V$ using the same notation used for sets.
@@
is also needed. It represents the total number of connections for a worker within $S$.
For a given $u\in V$ and $w\in\mathcal{W}$ we define the following sets:
@@
With this notation, we can easily formulate the conditions or constraints needed to define valid cycles.
To ensure that all workers $w\in\mathcal{W}$ cycles loop back to their assigned base $b_{w}\in B\subset V$ we use the bases constraints ($\mathcal{C}_{\text{B}}$) Equation 11:
@@
As some of the workers might remain inactive, we introduce new binary parameters ${\mathbf{Y}_{B}=\left\{y_{w}\in\{0,1\}\right\}_{w\in\mathcal{W}}}$.
Usually, it is simpler to set $y_{w}=1$ for $\forall w\in\mathcal{W}$ and simplify all the rest of the problem by eliminating any of the references to $\mathbf{Y}_{B}$. This however might yield sub-optimal solutions or even create contradictions. Consider the case where there are less tasks than active workers.
As all tasks must be completed by a compatible (remember that compatibility is encoded in the edges of the graph) and active worker exactly once in any of the available approaches, we impose the task completion constraints ($\mathcal{C}_{\text{T}}$) Equation 12 and Equation 13:
@@
We also introduce the parameters ${\mathbf{Y}_{C}=\{y_{v|w}\in\{0,1\}\}_{w\in\mathcal{W};\,v\in T}}$ to select which worker is to complete each of the tasks and define ${\mathbf{Y}=\mathbf{Y}_{C}\cup\mathbf{Y}_{B}}$. In a similar fashion to that of $\mathbf{Z}$, they are separated into fixed and free parameters. If a task $\tau$ is incompatible with $w$ ($\Gamma_{w}(\tau)=0$), then, $y_{v_{\tau}|w}$ is fixed at 00. All the parameters that correspond to compatible associations are left free to the solver. These constraints also enforce continuity between transitions.
Still, it is possible that the routes yielded by a specific selection of $\mathbf{Z}$ and ${\mathbf{Y}=\mathbf{Y}_{B}\cup\mathbf{Y}_{C}}$ contains several closed loops that do not share any common vertex. These are called subtours and they are an unwanted side effect that we need to remove from the model. To do so, we need to use some kind subtour elimination constraints ($\mathcal{C}_{\text{S}}$). In literature, two types are mainly used: the Dantzig–Fulkerson–Johnson [14] (DFJ SECs) or the Miller-Tucker-Zemlim [15] subtour elimination constraints (MTZ SECs). Each of them has their own unique pros and cons.
@@
$\left\{\Sigma^{(w)}S\leq|S|-1\right\}_{\begin{subarray}{c}W\in\mathcal{W}\\ S\in\{Q\subset V\,|\,b_{u}\notin Q,\,|Q|\geq 2\}\end{subarray}}$ (14)
Each of them enforces that for a given subset $Q\subset V$, the number of moves done by worker within the subset is at most the minimum number of moves needed to transverse all of its vertices. This way, no valid route can loop to itself within $Q$. As we know where each of the workers starts and finishes, we do not need to add the condition for all the possible subsets of $V$, just for those that do not contain the base.
The main advantage of these constraints is their simplicity and generality that allow for a more dynamic solving process without any modification to their original constraint’s formulation. The number of subtour constraints scale exponentially as so does the number of possible subsets $Q$. For instances where the number of tasked vertices is significant, the solving process with all the necessary constraints may take an absurd amount of computation effort so a different approach is needed. Instead of solving the complete problem, we relax it by removing the DFJ constraints. We then solve this relaxed instance and check for subtours. If subtours are found, we can eliminate them by adding the corresponding DFJ from Equation 14 for all $Q$ that are a subset of the set of vertices within each of the invalid subtours and then iterate again in a branch-and-cut fashion. In many cases, this approach speeds things up by orders of magnitude depending on the specific problem and alleviate some of the computing requirements. However, this heavily depends on the number of iterations one needs to make until the true optimal solution is found. It is also very limited in terms of extensions to the formulation. The DFJ SECs can also be used to forbid specific possible solutions to the problem.
On the other hand, although the process is not straightforward, the MTZ SECs can be easily adapted to this formulation by modifying the originals to the following:
$\left\{p_{u|w}-p_{v|w}+1\leq m_{w}(1-z_{[u,v]|w})\right\}_{\begin{subarray}{c}w\in\mathcal{W}\\ u,v\in T\lvert_{w}\end{subarray}}$ (15)
and by introducing two auxiliary constraints:
@@
$\left\{z_{[b_{w},u]|w}\leq p_{u|w}\right\}_{\begin{subarray}{c}w\in\mathcal{W}\\ u\in\delta^{(w)}_{+}(b_{w})\end{subarray}},$ (17)
where $m_{w}=|T\lvert_{w}|$ and $\left\{p_{u|w}\in\mathbb{N}\right\}_{w\in\mathcal{W};\,u\in T\lvert_{w}}=\mathbf{P}$ are new free parameters that encode the order at which a vertex $u\in T\lvert_{w}$ is visited within the route of $w\in\mathcal{W}$. The constraints Equation 15 enforces that if $u$ is visited before $v$ by $w$, then $p_{u|w}\geq p_{v|w}$. Notice however that if $v$ is visited immediately after $v$, $p_{v|w}$ might not be equal to $p_{u|w}+1$ as the number of vertices visited within a route might be less than $m_{w}$ so Equation 16 is trivially fulfilled. At the same time, Equation 16 and Equation 17 assigns $p=0$ for non-visited vertices and at least $p=1$ to the first one in each of the routes respectively. These constraints can be easily tightened so the values of $\mathbf{P}$ directly match with the visit order by defining $m_{w}=\sum_{e\in E\lvert_{w}}z_{e}$ and linearizing the products of the type $z_{e}m_{w}$. This however is not very useful and much of the actual useful characteristic of the MTZ SECs are also available in this formulation.
As the original MTZ SECs, they provide a natural framework for visit order constraints, however, they cannot be used for a relax-solve-iterate solving approach as they all must be included simultaneously to function properly. A mixed approach using both MTZ and DFJ might be feasible however.
Depending on the problem and in the case that the MTZ SECs are used, it might be useful to add constraints into the order of visit of each vertex in a route. For example, consider that the worker $w$ must visit $u$ before $v$ (denoted as $u<_{w}v$) for some technical reason. To implement this behavior, we would need to impose a new type of constraint in the model, the order of visit constraints $\mathcal{C}_{\text{O}}$. For each of the workers $w$ and pair of order-constrained vertices $[u,v]$, the following constraint must be added: